# Safe and Economical UAV Trajectory Planning in Low-Altitude Airspace: A Hybrid DRL-LLM Algorithm with Compliance Awareness

### v2

@@ Abstract
The rapid growth of the low-altitude economy has driven the widespread adoption of unmanned aerial vehicles (UAVs). This growing deployment presents new challenges for UAV trajectory planning in complex urban environments. However, existing studies often overlook key factors, such as urban airspace constraints and economic efficiency, which are essential in low-altitude economy contexts. Deep reinforcement learning (DRL) is regarded as a promising solution to these issues, while its practical adoption remains limited by low learning efficiency. To overcome this limitation, we propose a novel UAV trajectory planning algorithm that integrates DRL with the large language model (LLM) reasoning to enable safe, compliant, and economically viable trajectory planning. Specifically, we model the data collection task as a partially observable Markov decision process, explicitly incorporating obstacle avoidance, regulation awareness, and energy constraints. We design a hybrid optimization algorithm based on the soft actor-critic algorithm and LLM reasoning to enable adaptive decision-making in uncertain and dynamic environments. Experimental results demonstrate that our algorithm achieves the best overall performance, with the highest data collection rate (99.50%), almost zero collision avoidance rate and regulation violation rate, perfect successful landing rate (100%), and the lowest energy consumption rate (76.95%). These results validate the effectiveness of our algorithm in addressing UAV trajectory planning key challenges under constraints of the low-altitude economy networking.
@@ 1 Introduction
With the rapid advancement of the low-altitude economy, low-altitude airspace has emerged as a critical strategic resource to drive urban intelligence, industrial upgrading, and digital economic transformation [1]. This evolving domain integrates diverse sectors such as general aviation and urban air mobility, wherein unmanned aerial vehicles (UAVs) play an increasingly vital role in applications including logistics, agricultural monitoring, infrastructure inspection, and environmental protection [2]. Efficient utilization of this low-altitude airspace presents unprecedented opportunities for industry growth and technological innovation.
In this context, UAV-based data acquisition represents a fundamental technology enabling high-precision, multi-modal environmental sensing [3]. The UAV data acquisition process typically involves task assignment, flight trajectory planning, sensor data collection, and data transmission and processing. Effective trajectory planning is crucial to maximize area coverage and data quality, thereby supporting downstream applications with accurate and timely information [4].
@@
Challenge 1: Robust Obstacle Avoidance in Urban Airspaces. Due to the existence of airspace reuse in the low-altitude economy [36], operating in low-altitude urban environments exposes the data collection UAVs (DCU) not only to static obstacles (e.g., buildings) but also to dynamic obstacles (e.g., other UAVs). These obstacles are often partially observable and highly uncertain. Moreover, the DCU must also avoid restricted no-fly zones (NFZs) over sensitive areas. Ensuring robust obstacle avoidance under such conditions remains a core challenge. Traditional optimization methods typically lack the adaptability to handle such stochastic environments, and standard deep reinforcement learning (DRL) struggles to accumulate sufficient successful trajectories, leading to low learning efficiency [5].
@@
Research Gap. Most existing works have focused on specific aspects of UAV trajectory planning while neglecting others. In terms of traditional algorithms, [7] and [9]-[13] optimized communication and energy efficiency but did not consider obstacle avoidance. Although [8] and [14] incorporated obstacle avoidance, they were limited to static obstacles and overlooked dynamic threats like other UAVs. Meanwhile, DRL has been increasingly adopted for UAV-related optimization tasks due to its capacity to adaptively control in dynamic environments [18]. Studies using DRL, such as [27], focused on static obstacle avoidance or energy efficiency, yet lacked mechanisms for dynamic obstacle avoidance. Besides, [19]-[26] neglected critical aspects such as obstacle avoidance and adherence to regulatory constraints, limiting their applicability in real-world low-altitude environments. In parallel, large language models (LLMs) have also been introduced in mobile and aerial network scenarios to enhance intelligent decision-making and task coordination  [15]-[17]. Although some existing studies  [28]-[32] applied the LLM to UAV trajectory planning, they primarily focused on obstacle avoidance while neglecting compliance requirements and the need for data collection.
In summary, existing methods either focus on isolated objectives or assume simplified environments, and thus fail to address all the challenges mentioned above of the DCU trajectory planning. To address these limitations, this paper proposes a novel hybrid DRL and the LLM algorithm for safe and economical DCU trajectory planning in low-altitude airspace with compliance-awareness. Specifically, we introduce an LLM to enhance the DCU’s decision-making capability in complex, uncertain scenarios where rule compliance, contextual awareness, and obstacle interpretation are critical. While the LLM offers strong reasoning abilities and can infer safe actions based on high-level semantic input, it lacks the ability to learn optimal control strategies through interaction with dynamic environments [33]. Conversely, DRL excels at learning adaptive control policies via trial and error, but struggles to incorporate abstract constraints such as regulatory rules [34]. Therefore, we combine DRL with LLM to exploit their complementary strengths. The main contributions are as follows:
  • Mathematical modeling of DCU trajectory planning problem for low-altitude data collection. We model the DCU trajectory planning for low-altitude data acquisition task as a mathematical problem. This problem captures the robust obstacle avoidance requirement under both static and dynamic obstacles mentioned in Challenge 1.
  • Problem formulation as a partially observable Markov decision process (POMDP). We formulate the mathematical problem as a POMDP to account for partial observability in the low-altitude airspace. To capture specific requirements, such as the robust obstacle avoidance mentioned in Challenge 1 and the requirement of economy and compliance mentioned in Challenge 2, we carefully design the state space, action space, and reward function.
@@
Extensive experiments demonstrate that our proposed algorithm outperforms baseline approaches. Compared to the best-performing baseline, our algorithm improves the data collection rate by 2.9%, reduces the collision rate and regulation violation rate to almost 0 %, maintains 100% success landing rate, and lowers the energy consumption rate by 1.9%, thereby effectively addressing the above three challenges.
@@ 2 Related Work
In this section, we provide related works on planning UAV trajectories using traditional algorithms, DRL, and the LLM. The summary of related work is shown in Table I.
@@ 2.1 UAV Trajectory Planning with Traditional Algorithms
Qin et al. [7] employed a joint optimization algorithm using Dinkelbach’s method and block coordinate descent (BCD) to optimize trajectory, power, bit allocation, and phase shift in reconfigurable intelligent surface (RIS)-assisted UAV-enabled mobile edge computing (MEC) systems. Pan et al. [8] proposed a joint UAV power and trajectory planning framework using an improved non-dominated sorting genetic algorithm-II with K-means initialization operator and variable dimension mechanism (NSGA-II-KV) and an improved particle swarm optimization with normal distribution initialization, genetic mechanism, differential mechanism, and pursuit operator (PSO-NGDP). Pervez et al. [9] studied a multi-UAV MEC system and applied game theory and successive convex approximation (SCA) for joint optimization. Zhang et al. [10] used an elite-driven differential evolution algorithm with dynamic population evolution to jointly optimize UAV deployment and flight planning in Internet of Things (IoT) data collection. Heo et al. [11] incorporated an NFZ avoidance constraint into UAV trajectory planning and used quadratic transform and SCA for joint optimization. Li et al. [12] applied a swarm intelligence-based algorithm (SIA) to a multi-objective problem involving collaborative beamforming for energy-efficient data harvesting. Fu et al. [13] adopted the distributed approximate Newton (DANE) method and BCD to optimize UAV trajectory, device scheduling, and resource allocation. Wang et al. [14] utilized an ant colony optimization (ACO)-based algorithm for global obstacle avoidance and communication in RIS-assisted UAV systems.
@@ 2.2 UAV Trajectory Planning with DRL
Silvirianti et al. [19] proposed a layerwise quantum-based DRL (LQ-DRL) method for UAV trajectory planning, user grouping, and power allocation. Chen et al. [20] formulated a multi-stage mixed-integer nonlinear programming problem and proposed the joint trajectory planning and resource allocation by DRL for UAV-assisted MEC. Ning et al. [21] addressed distributed trajectory control using DRL and game theory in multi-UAV MEC networks. Song et al. [22] modeled a multi-objective Markov decision process with a trace-based experience replay scheme for energy-efficient UAV trajectory planning. Ding et al. [23] proposed multi-agent advantage actor-critic with deep deterministic policy gradient (DDPG) and deep attention-based inference (DAI) schemes for spectrum allocation and UAV trajectory design. Wang et al [24] developed a safe DRL algorithm based on a constrained Markov decision process for UAV trajectory planning. He et al. [25] proposed a distributed multi-objective trajectory design method using distributed DRL for trajectory planning of multi-UAV. Liu et al. [26] applied multi-agent RL (MARL) and SCA to optimize UAV trajectory in a cell-free space-air-ground integrated network. Ning et al. [27] used constrained DRL to optimize UAV movement and connection assignment in wireless-powered IoT networks.
@@ 2.3 UAV Trajectory Planning with LLM
Zhong et al. [28] presented a vision-based autonomous planning system for UAVs in dynamic environments. It combined lightweight object detection and Kalman filtering to track moving obstacles. Then they explored integrating an LLM for more natural human-UAV interaction. Phadke et al. [29] explored the integration of the LLM into UAV control, enabling natural language interaction for more intuitive and accessible operations. It highlighted potential benefits in decision-making, mission planning, and safety, and presented a development framework with proof-of-concept demonstrations using existing LLMs and simulation platforms. Xiao et al. [30] introduced a foundation model-guided trajectory planning framework that integrated LLMs and vision-language models (VLMs) for enhanced semantic reasoning and visual perception in UAV navigation. Samma et al [31] proposed a novel LLM-based visual trajectory planning method for UAVs operating in indoor GPS-denied environments. By interpreting camera data on detected persons and depth, the fine-tuned LLM generated safe and efficient flight paths. Cai et al. [32] presented a hybrid autonomous UAV landing system that combined LLM with model predictive control (MPC) to improve safety and precision in dynamic, unstructured environments.
@@ 2.4 Discussion
Although numerous studies have explored UAV trajectory planning for applications such as data collection, obstacle avoidance remains an under-addressed challenge. For instance, while works [7]-[10] focused on joint optimization of resources and trajectories, they either entirely overlooked obstacle avoidance or considered only static obstacles, failing to avoid dynamic obstacles in the low-altitude airspace. Although [11] incorporated NFZ, it neglected other essential regulatory constraints. Similarly, [12]-[14] did not integrate any obstacle avoidance mechanisms, or again, only addressed static obstacles. Studies [19]-[23] ignored obstacle avoidance altogether, and although [24] introduced learning-based safety mechanisms, it did not jointly consider energy constraints and obstacle avoidance. Furthermore, [25] and [26] omitted obstacle considerations, while [27] did not account for dynamic obstacles such as other UAVs. Additionally, although studies [28]-[30] employed an LLM for UAV trajectory planning, they only considered the avoidance of a single obstacle in the planning process. While [31] and [32] offered a more comprehensive approach to obstacle avoidance, they did not account for compliance, low resource consumption, or data collection requirements.
To overcome these issues, this paper introduces an innovative algorithm that integrates DRL with the LLM to enable safe, compliance-aware, and cost-effective UAV trajectory planning in low-altitude airspace.
@@
Pan et al. 2024 [8] NSGA-II-KV, PSO-NGDP Static only
@@
Li et al. 2024 [12] SIA
Fu et al. 2025 [13] DANE, BCD
Wang et al. 2025 [14] ACO Static only
@@
Figure 1: System model of low-altitude UAV data collection. The DCU collects data from GE while avoiding static obstacles (BZs, NFZs, RZs) and dynamic obstacles (OUs) under energy and compliance constraints.
@@
${{\mathbb{R}}_{XX,x}}=({{p}_{XX,x}},$ ${{l}_{XX,x}},{{w}_{XX,x}})$ Location of the $x$-th XX in $\mathbb{A}$ (XX = NFZ, BZ, or RZ)
${N}_{XX}$} Number of XX (XX = GE, OU, NFZ, BZ, or RZ)
${{L}_{GE,i}}=({{x}_{GE,i}},$ ${{y}_{GE,i}},0)$ Location of the $i$-th GE
$D{{V}_{G{{E}_{i}},t}}$ Data volume of the $i$-th GE at time $t$
$T{{P}_{GE,i}}$ Transmit power of the $i$-th GE
${{L}_{XX,t}}=({{x}_{XX,t}},$ ${{y}_{XX,t}},{{z}_{XX,t}})$ Location of XX at time $t$ (XX = DCU or OU)
$SL=({{x}_{SL}},{{y}_{SL}},{{z}_{SL}})$ Starting location of the DCU
$EL=({{x}_{EL}},{{y}_{EL}},{{z}_{EL}})$ Ending location of the DCU
$H$ Fixed flight altitude of the DCU and OUs
PRPR Perception radius of the DCU
@@
${{v}_{XX}}(t)=({{v}_{XX}}{{(t)}_{x}},$ ${{v}_{XX}}{{(t)}_{y}},{{v}_{XX}}{{(t)}_{z}})$ Flight speed of XX at time $t$ (XX = DCU or OU)
${v}_{\text{limit}}$} Speed limit in ${\mathbb{R}}_{RZ}$} for the DCU
${E}_{total}$} Total energy the DCU has
${E}_{\text{limit}}$} Maximum energy that the DCU can consume
$E(t)$ Energy consumed during time slot $\Delta t$
$e(t)$ Remaining energy at time $t$
${{e}_{fly}}(t)$ Flight energy consumption during $\Delta t$ starting at $t$
${{e}_{hover}}(t)$ Hover energy consumption during $\Delta t$ starting at $t$
@@
$T$ Total time of the data collection
${t}_{fly}$} Time used for flight in $\Delta t$
${t}_{col}$} Time used for data collection in $\Delta t$
$T{{D}_{u,G{{E}_{i}}}}(t)$ Data transmitted between DCU and GEi during $\Delta t$ starting at $t$
${{B}_{u,G{{E}_{i}}}}(t)$ Effective rate between DCU and GEi during $\Delta t$
$\wedge$ Logical AND
$\odot$ Hadamard product

### v3

@@ Abstract
The rapid growth of the low-altitude economy has driven the widespread adoption of unmanned aerial vehicles (UAVs). This growing deployment presents new challenges for UAV trajectory planning in complex urban environments. However, existing studies often overlook key factors, such as urban airspace constraints and economic efficiency, which are essential in low-altitude economy contexts. Deep reinforcement learning (DRL) is regarded as a promising solution to these issues, while its practical adoption remains limited by low learning efficiency. To overcome this limitation, we propose a novel UAV trajectory planning algorithm that integrates DRL with the large language model (LLM) reasoning to enable safe, compliant, and economically viable trajectory planning. Specifically, we model the trajectory planning task as a partially observable Markov decision process, explicitly incorporating obstacle avoidance, regulation awareness, and energy constraints. We design a hybrid optimization algorithm based on the soft actor-critic algorithm and LLM reasoning to enable adaptive decision-making in uncertain and dynamic environments. Experimental results demonstrate that our algorithm achieves the best overall performance, with the highest data collection rate (99.50%), almost zero collision avoidance rate and regulation violation rate, perfect successful landing rate (100%), and the lowest energy consumption rate (76.95%). These results validate the effectiveness of our algorithm in addressing UAV trajectory planning key challenges under constraints of the low-altitude economy networking.
@@ 1 Introduction
With the rapid advancement of the low-altitude economy, low-altitude airspace has emerged as a critical strategic resource to drive urban intelligence, industrial upgrading, and digital economic transformation [1, 2]. This evolving domain integrates diverse sectors such as general aviation and urban air mobility, wherein unmanned aerial vehicles (UAVs) play an increasingly vital role in applications including logistics, agricultural monitoring, infrastructure inspection, and environmental protection [3]. However, the high-density and regulation-intensive nature of low-altitude airspace also poses unique challenges for intelligent trajectory planning and safe autonomous operation.
In this context, UAV-based data acquisition represents a fundamental technology enabling high-precision, multi-modal environmental sensing [4]. The UAV data acquisition process typically involves task assignment, flight trajectory planning, sensor data collection, and data transmission and processing. Effective trajectory planning is crucial to maximize area coverage and data quality, thereby supporting downstream applications with accurate and timely information [5]. It also serves as a key enabler for intelligent and autonomous decision-making in dense and dynamic airspaces.
@@
Challenge 1: Robust Obstacle Avoidance in Urban Airspaces. Due to the existence of airspace reuse in the low-altitude economy [32], operating in low-altitude urban environments exposes the data collection UAVs (DCU) not only to static obstacles (e.g., buildings) but also to dynamic obstacles (e.g., other UAVs). These obstacles are often partially observable and highly uncertain. Moreover, the DCU must also avoid restricted no-fly zones (NFZs) over sensitive areas. Ensuring robust obstacle avoidance under such conditions remains a core challenge. Ensuring robust obstacle avoidance under such conditions remains a core challenge. Traditional optimization methods typically lack the adaptability to handle such stochastic environments, and standard deep reinforcement learning (DRL) struggles to accumulate sufficient successful trajectories, leading to low learning efficiency [6].
@@
Most existing methods address only isolated objectives or assume simplified environments, which prevents them from achieving a unified balance among safety, compliance, and economy. To address these limitations, this paper proposes a novel hybrid DRL and large language model (LLM) algorithm for safe and economical DCU trajectory planning in low-altitude airspace with compliance-awareness. The motivation is that, while the LLM offers strong reasoning abilities and can infer safe actions based on high-level semantic input, it lacks the ability to learn optimal control policies through interaction with dynamic environments [30]. Conversely, DRL excels at learning adaptive control policies via trial and error, but struggles to incorporate abstract constraints such as regulatory rules [31]. Therefore, we introduce an LLM to enhance the DCU’s decision-making capability in complex, uncertain scenarios where rule compliance, contextual awareness, and obstacle interpretation are critical. The main contributions of this paper are summarized as follows:
  • Unified modeling of low-altitude UAV trajectory planning. We formulate the DCU trajectory planning task for low-altitude data acquisition as a partially observable Markov decision process (POMDP), which jointly captures constraints including both static and dynamic obstacle avoidance mentioned in Challenge 1, regulation awareness, and energy efficiency mentioned in Challenge 2, under partial observability.
@@
  • Comprehensive experimental validation and performance improvement. Extensive experiments verify that the proposed algorithm consistently outperforms state-of-the-art baselines across all key metrics. Compared with the best-performing baseline, our method improves the data collection rate by 2.9%, reduces the collision rate and regulation violation rate to nearly 0%, achieves a 100% successful landing rate, and lowers the energy consumption rate by 1.9%. These results confirm the algorithm’s effectiveness in addressing the Challenges 1-3.
@@ 2 Related Work
In this section, we review existing UAV trajectory planning methods from three perspectives, including traditional optimization algorithms, DRL-based approaches, and LLM-enhanced methods. Table I summarizes representative studies and their characteristics.
@@ 2.1 UAV Trajectory Planning with Traditional Algorithms
Early studies mainly adopted optimization-based algorithms to jointly optimize UAV trajectory, communication, and energy efficiency. Qin et al. [8] used Dinkelbach’s method and block coordinate descent (BCD) for joint trajectory and resource optimization in reconfigurable intelligent surface-assisted UAV mobile edge computing (MEC) systems. Pan et al. [9] combined an improved non-dominated sorting genetic algorithm II (NSGA-II) with a customized particle swarm optimization (PSO) variant for multi-objective power and trajectory planning. Pervez et al. [10] formulated a multi-UAV MEC optimization using game theory and successive convex approximation (SCA). Zhang et al. [11] applied differential evolution to balance UAV deployment and flight planning for Internet of Things (IoT) data collection. Heo et al. [12] incorporated NFZ constraints using quadratic transform and SCA. Other works [13, 14, 15] explored swarm intelligence and distributed optimization for beamforming and energy efficiency.
Traditional optimization algorithms are effective in structured or static environments but lack adaptability to dynamic and uncertain low-altitude airspace conditions, and they cannot efficiently handle the stochastic nature of real-world obstacle distributions (Challenge 1).
@@ 2.2 UAV Trajectory Planning with DRL
To improve adaptability, many studies employed DRL for UAV trajectory control and resource management. Silvirianti et al. [16] proposed a layerwise quantum DRL method for joint trajectory and power optimization. Chen et al. [17] applied DRL to solve a multi-stage mixed-integer problem for UAV-assisted MEC. Ning et al. [18] introduced distributed DRL with game theory for collaborative control. Song et al. [19] formulated a multi-objective Markov decision process (MDP) with a trace-based experience replay mechanism. Ding et al. [20] developed a multi-agent advantage actor-critic (A2C)–deep deterministic policy gradient (DDPG) scheme with attention-based inference. Wang et al. [21] incorporated safety constraints into DRL under a constrained MDP framework. He et al. [22] and Liu et al. [23] extended multi-agent DRL to cooperative trajectory planning. Ning et al. [24] adopted constrained DRL for UAV mobility and connection optimization in IoT networks.
DRL-based methods can adapt to partially observable and dynamic environments, but they often require extensive exploration to incorporate domain-specific rules such as regulatory or economic constraints (Challenge 2 and Challenge 3).
@@ 2.3 UAV Trajectory Planning with LLM
Recent research has introduced LLMs to enhance UAV autonomy by integrating reasoning and natural language understanding. Zhong et al. [25] combined lightweight object detection with an LLM for vision-based UAV planning in dynamic scenes. Phadke et al. [26] explored LLM-driven UAV control via natural language interfaces for improved mission coordination and safety. Xiao et al. [27] proposed a foundation model-guided trajectory framework integrating LLMs and vision-language models (VLMs) for semantic reasoning and perception. Samma et al. [28] utilized an LLM-based vision system to plan UAV trajectories in GPS-denied indoor environments. Cai et al. [29] designed an LLM–model predictive control (MPC) hybrid landing system for enhanced safety in unstructured scenarios.
LLM-enhanced approaches offer semantic reasoning and interpretability but face challenges in real-time decision-making for UAVs, including excessive resource consumption and limited real-time performance.
@@ 2.4 Discussion
Existing UAV trajectory planning methods contribute valuable insights but exhibit limitations when applied to low-altitude data collection missions. Traditional algorithms [8]-[15] cannot adapt to dynamic and uncertain obstacle patterns in real urban airspaces. DRL-based methods [16]-[24] improve adaptability but overlook regulatory compliance and economic trade-offs. LLM-based frameworks [25]-[29] offer strong reasoning but lack real-time adaptability. Consequently, none of these approaches can simultaneously ensure robust obstacle avoidance, compliance with evolving regulations, and energy-efficient trajectory design in low-altitude environments. To overcome these issues, this paper proposes a hybrid DRL–LLM algorithm that integrates adaptive control and semantic reasoning to achieve safe, compliant, and economical UAV trajectory planning.
@@
Pan et al. 2024 [9] NSGA-II, PSO variant Static only
@@
Li et al. 2024 [13] Swarm intelligence-based algorithm
Fu et al. 2025 [14] Distributed approximate Newton, BCD
Wang et al. 2025 [15] Ant colony optimization Static only
@@
Figure 1: System model of low-altitude UAV trajectory planning. The DCU collects data from GE while avoiding static obstacles (BZs, NFZs, RZs) and dynamic obstacles (OUs) under energy and compliance constraints.
@@
${B}_{\mathrm{GE}_{i}}(t)$ Effective rate between the DCU and GEi in the time slot $t$
$d_{t,\mathrm{XX}}$ Distance between the DCU and XX at the beginning of time slot $t$ (XX = $\mathbb{A}$, $\mathbb{LA}$, OU, BZ, or NFZ)
$d_{\mathrm{min},\mathrm{XX}}$ Minimum distance between the DCU and XX (XX = $\mathbb{A}$, OU, BZ, or NFZ)
$d_{\mathrm{safe},\mathrm{XX}}$ Safe distance between the DCU and XX (XX = $\mathbb{A}$, OU, BZ, or NFZ)
$D{{V}_{G{{E}_{i}},t}}$ Data volume of the $i$-th GE at the beginning of time slot $t$
$E(t)$ Remaining energy at the beginning of time slot $t$
${E}_{\text{limit}}$} Maximum energy that the DCU can consume
${E}_{\mathrm{total}}$} Total energy the DCU has
$e(t)$ Energy consumed in the time slot $t$ of the DCU
$e_{\mathrm{fly}}(t)$ Flight energy consumption in the time slot $t$ of the DCU
$e_{\mathrm{hover}}(t)$ Hover energy consumption in the time slot $t$ of the DCU
$\mathrm{EL}=({{x}_{\mathrm{EL}}},{{y}_{\mathrm{EL}}},{{z}_{\mathrm{EL}}})$ Ending location of the DCU
$H$ Fixed flight altitude of the DCU and OUs
${{L}_{{\mathrm{GE}}_{i}}}=({{x}_{{\mathrm{GE}}_{i}}},{{y}_{{\mathrm{GE}}_{i}}},0)$ Location of the $i$-th GE
${{L}_{\mathrm{XX},t}}=({{x}_{\mathrm{XX},t}},{{y}_{\mathrm{XX},t}},{{z}_{\mathrm{XX},t}})$ Location of XX at the beginning of time slot $t$ (XX = DCU or OU)
${N}_{XX}$} Number of XX (XX = GE, OU, NFZ, BZ, or RZ)
PRPR Perception radius of the DCU
$\mathrm{SL}=({{x}_{\mathrm{SL}}},{{y}_{\mathrm{SL}}},{{z}_{\mathrm{SL}}})$ Starting location of the DCU
${\mathrm{SNR}}_{\mathrm{GE}_{i}}(t)$ The signal-to-noise ratio at the beginning of time slot $t$ of $\mathrm{GE}_{i}$
@@
$T$ Total number of time slots
$TD_{{\mathrm{GE}}_{i}}(t)$ Data transmitted between the DCU and GEi in the time slot $t$
$T{{P}_{{\mathrm{GE}}_{i}}}$ Transmit power of the $i$-th GE
${t}_{\mathrm{fly}}$} Time used for flight in the time slot $t$ of the DCU
${t}_{\mathrm{hover}}$} Time used for hovering in the time slot $t$ of the DCU
${v}_{\mathrm{DCU}}(t)=({{v}_{x}}(t),{v}_{y}(t),{{v}_{z}}(t))$ Flight speed of the DCU at the beginning of time slot $t$
@@
${v}_{\text{limit}}$} Speed limit in ${\mathbb{R}}_{\mathrm{RZ}}$} for the DCU and OUs
$\wedge$ Logical AND operator
$\odot$ Hadamard (element-wise) product
${{\mathbb{R}}_{\mathrm{XX}_{x}}}=({{p}_{\mathrm{XX}_{x}}},{{l}_{\mathrm{XX}_{x}}},{{w}_{\mathrm{XX}_{x}}})$ Location of the $x$-th XX in $\mathbb{A}$ (XX = NFZ, BZ, or RZ)
@@
$\tau$ The threshold of ${\mathrm{SNR}}_{\mathrm{GE}_{i}}(t)$